# Light Edge Fault Tolerant Graph Spanners

### v1

@@ 2.1 Overview of Upper Bounds
Next, we explain the relationship to the weighted girth framework from [ENS14]. Standard arguments (first introduced by [BP19]) can be adapted to our modified algorithm to show that essentially every cycle in $H$ with at most $k$ edges must contain two edges that form a block, i.e., the blocks cover all cycles with at most $k+1$ edges. Following a proof from [ENS14], this argument can be adapted to show that they also cover essentially all cycles with “normalized weight” at most $k+1$, i.e., cycles $C$ where $w(C)/\max_{e\in C}w(e)$ is at most $k+1$. This is useful because if there were no cycles of normalized weight at most $k+1$ — or, in the language of [ENS14], $H$ has weighted girth $>k+1$ — then by definition $H$ has lightness at most $\lambda(n,k+1)$. So our goal will be to turn our spanner $H$ where low-normalized-weight cycles are blocked into subgraph(s) $H^{\prime}$ where there are no low-normalized-weight cycles, in a way that allows us to bound the lightness of $H$ using the lightness of $H^{\prime}$.
@@
In the sparsity setting, one of the now-standard ways of handling a blocking set is by subsampling the graph, showing that in the resulting graph no blocks survive (and thus there cannot be any short cycles and so there are not many edges), and then using this to argue that in the original graph there cannot be many edges (since we just subsampled with a known probability). A natural first attempt is to use the same idea for lightness, but it immediately runs into a serious problem. The connectivity preserver $Q$ itself can have unblocked cycles of low normalized weight, and so if the subsampling does not remove any edges of $Q$ then we cannot hope to get rid of all such cycles. However, if the subsampling does get rid of edges of $Q$, then the min-weight connectivity preserver of the subsampled graph $H^{\prime}$ might be very different from $Q$, making it hard to compare the lightnesses of $H$ and $H^{\prime}$. (Note that this problem does not show up when analyzing sparsity, since the number of edges is not normalized by anything and so it is relatively easy to compare $|E(H)|$ to $|E(H^{\prime})|$).
@@ 3.2 The Greedy Algorithm and Blocking Sets
The output spanner $H$ from the fault-tolerant greedy algorithm has an $f$-capped edge-blocking set $B$ that blocks all cycles $C$ that contain at least one edge from $G\setminus Q$ and which have normalized weight $w^{*}(C)\leq k+1$. Additionally, for every pair $(e,e^{\prime})\in B$, the first edge $e$ is not in $Q$.
@@
Since each $|F_{e}|\leq f$, we have that $B$ is $f$-capped. To argue its cycle-blocking properties, let $C$ be a cycle as in the lemma statement, and let $(u,v)\in C$ be the last edge in $C$ to be considered in the main loop of the algorithm. Just before $(u,v)$ is added to $H$, there exists a $u\leadsto v$ path through the other edges of the cycle, and so we have
@@ 3.3 Warmup: Slow Suboptimal Competitive Lightness Upper Bounds
Our next step is to construct a sequence of three subgraphs $H[T]\supseteq H^{\prime}[T]\supseteq H^{\prime\prime}[T]$ associated to each tree $T\in\mathcal{T}$. In the following let $B$ be an $f$-capped blocking set for $H$ as in Lemma 20.
@@
  • (Construction of $H^{\prime\prime}[T]$) Finally, to construct $H^{\prime\prime}[T]$ from $H^{\prime}[T]$, for every pair $(e,e^{\prime})\in B$ with $e,e^{\prime}\in H^{\prime}[T]$, we delete the first edge $e$, and let $H^{\prime\prime}[T]$ be the remaining subgraph.
@@
Lemma 26 (Subgraph Properties).
Every graph $H^{\prime\prime}[T]$ has weighted girth $>k+1$ (deterministically), and expected weight
@@
First, let $C$ be any cycle in $H[T]$ of normalized weight $\leq k+1$. We will argue that $C$ does not survive in $H^{\prime\prime}[T]$. Note that $C$ contains at least one edge from $E(H)\setminus Q$, since the only edges from $Q$ in $H[T]$ are those in $T$ itself, which is a tree. Thus, by the properties of the blocking set $B$ from Lemma 20, there is $(e,e^{\prime})\in B$ with $e,e^{\prime}\in C$ and $e\notin T$. It is not possible for both edges $e,e^{\prime}$ to survive in $H^{\prime\prime}[T]$, since if they are both selected to remain in $H^{\prime}[T]$ then we will choose to delete $e$ when we construct $H^{\prime\prime}[T]$. Thus $C$ does not survive in $H^{\prime\prime}[T]$.
Next, we measure the expected weight in $H^{\prime\prime}[T]$. Every edge in $T$ is included in $H^{\prime\prime}[T]$ deterministically. Every edge $e\in E(H[T])\setminus T$ survives in $H^{\prime\prime}[T]$ iff the following two independent events both occur:
@@
Thus all edges $e\in E(H[T])$ survive in $H^{\prime\prime}[T]$ with probability $\Omega(1/f)$, implying the lemma. ∎
@@
As in the probabilistic method, there exists a realization of the subgraphs $H^{\prime\prime}[T]$ that satisfies the expected weight inequality in the previous lemma (for all $T$). Using this, we observe that for all trees $T\in\mathcal{T}$, we have
$\frac{w(H[T])}{w(T)}$ $\leq\frac{O\left(f\cdot w(H^{\prime\prime}[T])\right)}{w(T)}\leq O(f)\cdot\frac{w(H^{\prime\prime}[T])}{w(\texttt{mst}(H^{\prime\prime}[T]))}=O(f)\cdot\ell(H^{\prime\prime}[T])$
$=O(f)\cdot\lambda(n,k+1)$ (2)
where the last inequality follows from the definition of $\lambda$ and the previous lemma establishing that $H^{\prime\prime}[T]$ has weighted girth $>k+1$. Thus, to wrap up the proof, we bound:
@@ 3.4 Improved Lightness Bounds via Multiple Host Trees
One way in which our previous analyses are suboptimal is that they do not acknowledge the possibility that an edge could be hosted by many trees, not just one. As a simple example to introduce the technique, let us observe that the lightness bound from Algorithm 1 improves by a factor of $f$ if we are willing to take a higher lightness competition parameter:
@@
Follow the previous analysis, but note that our guarantee is now that we have a set of trees $\mathcal{T}$ such that (1) every edge is in at most two trees, and (2) for all $(u,v)\in E(G)\setminus Q$, there are $(2+\eta)f+1$ trees in $\mathcal{T}$ that span the $(2+\eta)f+1$-connected component that contains $u,v$. Thus, letting $F$ be a fault set that caused us to add $(u,v)$ to the spanner, there are $\Theta(\eta f)$ of these trees that are disjoint from $F$. We let each of these trees host the edge $(u,v)$. In particular, this means that $(u,v)$ will appear in $\Theta(f)$ many $H[T]$ subgraphs. From there the analysis continues as before, but our final calculation is:
@@ 3.5 Competitive Lightness Upper Bounds in Polynomial Time
The Chernoff bound (see e.g. [DP09] for exposition) states that:
@@
The following lemmas will assume that the event from Lemma 29 holds, and hence these lemmas only with high probability, rather than deterministically. We next establish correctness:
@@
There are $|F|\leq f$ edges, which are each included in $H^{\prime}$ with probability $1/f$, and so none of the edges in $F$ survive in $H^{\prime}$ with probability at least $1/2$. So we have $P_{e}^{T}\geq 1/2$. Assuming that the event from Lemma 29 holds (relating $P^{T}_{e}$ to $\widehat{P}_{e}^{T}$), we thus have $\widehat{P}_{e}^{T}\geq 3/8$. So $(u,v)$ will be added to the spanner $H$ when the tree $T$ is considered, completing the proof. ∎
@@
Analogous to the previous warmup, it will be helpful to define subgraphs $H^{\prime\prime}[T]\subseteq H^{\prime}[T]\subseteq H[T]$ as follows. First, $H^{\prime}[T]$ is a random subgraph of $H[T]$, obtained by including $T$ deterministically and then including each edge hosted by $T$ with probability $1/f$. For an edge $(u,v)\in H^{\prime}[T]$, let $H^{\prime}_{(u,v)}[T]$ be the subgraph that contains only $T$ and the edges from $H^{\prime}[T]$ considered strictly before $(u,v)$ in the algorithm. Then we define $H^{\prime\prime}[T]$ as:
@@
We can now bound the expected weight of $H^{\prime\prime}[T]$ in two different ways:
  • On one hand, the definition of $H^{\prime\prime}[T]$ implies that it has weighted girth $>k+1$. To see this, consider a cycle $C$ in $H^{\prime}[T]$ of normalized weight $\leq k+1$ and let $(u,v)\in C$ be its last edge considered in the algorithm. If all edges in $C\setminus\{(u,v)\}$ are added to $H^{\prime\prime}[T]$, then they form a $u\leadsto v$ path, implying that
@@
    which will cause us to not include $(u,v)$ in $H^{\prime\prime}_{(u,v)}[T]$. Thus, by definition of $\lambda$, we have
    $\lambda(n,k+1)\geq\ell(H^{\prime\prime}[T])$ $=\frac{w(H^{\prime\prime}[T])}{w(\texttt{mst}(H^{\prime\prime}[T]))}$
    $\geq\frac{w(H^{\prime\prime}[T])}{w(T)}.$
@@
$w(H[T])$ $\leq O\left(f\cdot\mathbb{E}\left[w(H^{\prime\prime}[T])\right]\right)$
@@
The first part of Theorem 13 is now directly implied by Lemmas 30 and 32.
@@ 5 Future Directions: Light Vertex Fault Tolerant Spanners
In addition to the obvious open question of determineg the precise bound on the achievable $2f$-competitive lightness, an interesting open problem left by this paper is to determine the appropriate competition parameter for light vertex fault tolerant (VFT) spanners. We can define vertex competitive lightness analogously to Definition 8, but with respect to the set of $f$-VFT connectivity preservers rather than $f$-EFT connectivity preservers. We then ask:
Question 34.
What is the smallest function $\mathcal{C}(f)$ such that every $n$-node weighted graph $G$ has an $f$-VFT $k$-spanner with vertex-competitive lightness [^9] It would also be interesting if, instead of a $\lambda$ reduction, one could match the state-of-the art non-faulty size/stretch tradeoff up to a $\texttt{poly}(f)$ factor. That is, the goal would be to obtain stretch $(1+\varepsilon)(2k-1)$ and size $O\left(\texttt{poly}(f,\varepsilon^{-1})\cdot n^{1+1/k}\right).$

### v2

@@ 2.1 Overview of Upper Bounds
Next, we explain the relationship to the weighted girth framework from [ENS14]. Standard arguments (first introduced by [BP19]) can be adapted to our modified algorithm to show that essentially every cycle in $H$ with at most $k+1$ edges must contain two edges that form a block, i.e., the blocks cover all cycles with at most $k+1$ edges. Following a proof from [ENS14], this argument can be adapted to show that they also cover essentially all cycles with “normalized weight” at most $k+1$, i.e., cycles $C$ where $w(C)/\max_{e\in C}w(e)$ is at most $k+1$. This is useful because if there were no cycles of normalized weight at most $k+1$ — or, in the language of [ENS14], $H$ has weighted girth $>k+1$ — then by definition $H$ has lightness at most $\lambda(n,k+1)$. So we have a spanner $H$ where cycles of normalized weight at most $k+1$ exist but are blocked, and we know that if there were no such cycles then we have good lightness bounds. So our goal will be to turn $H$ into some other graph $H^{\prime}$ without any cycles of normalized weight at most $k+1$, in a way that allows us to bound the lightness of $H$ using the lightness of $H^{\prime}$.
@@
In the sparsity setting, one of the now-standard ways of going from a graph $H$ where short cycles exist but are blocked to a graph $H^{\prime}$ where there are no short cycles is by subsampling $H$ to get $H^{\prime}$. Informally, by choosing the correct probability, one can often show that no blocks survive in $H^{\prime}$, and thus there cannot be any short cycles and so there are not many edges in $H^{\prime}$. On the other hand, since we obtained $H^{\prime}$ by subsampling $H$ with a known probability, there are not too many more edges in $H$ than in $H^{\prime}$. Thus $H$ cannot have many edges. A natural first attempt is to use the same idea for lightness, but it immediately runs into a serious problem. The connectivity preserver $Q$ itself can have unblocked cycles of low normalized weight, and so if the subsampling does not remove any edges of $Q$ then we cannot hope to get rid of all such cycles. However, if the subsampling does get rid of edges of $Q$, then the min-weight connectivity preserver of the subsampled graph $H^{\prime}$ might be very different from $Q$, making it hard to compare the lightnesses of $H$ and $H^{\prime}$. (Note that this problem does not show up when analyzing sparsity, since the number of edges is not normalized by anything and so it is relatively easy to compare $|E(H)|$ to $|E(H^{\prime})|$).
@@ 3.2 The Greedy Algorithm and Blocking Sets
The output spanner $H$ from the fault-tolerant greedy algorithm has an $f$-capped edge-blocking set $B$ that blocks all cycles $C$ with normalized weight $w^{*}(C)\leq k+1$ and where a heaviest edge in $C$ is not in $Q$. Additionally, for every pair $(e,e^{\prime})\in B$, the first edge $e$ is not in $Q$.
@@
Since each $|F_{e}|\leq f$, we have that $B$ is $f$-capped. To argue its cycle-blocking properties, let $C$ be a cycle as in the lemma statement, with a heaviest edge $(u,v)\notin Q$. When we consider $(u,v)$ in the main loop of the algorithm, just before $(u,v)$ is added to $H$, there exists a $u\leadsto v$ path through the other edges of the cycle, and so
@@
A subtlety here is that the output spanner $H$ could still have unblocked cycles $C$ of low weighted girth, in the case where the heaviest edge of $C$ is in $Q$ (and so it was added before the main greedy loop of the algorithm), and the edges in $C\setminus Q$ are much lighter. In general, min-weight $f$-EFT connectivity preservers are less well-behaved than min-weight spanning trees, and so situations like this are indeed possible even despite $Q$ being min-weight. We will need to handle these cycles in our analysis in another way (see Lemma 26).
@@ 3.3 Warmup: Slow Suboptimal Competitive Lightness Upper Bounds
Our next step is to construct a sequence of four subgraphs $H[T]\supseteq H^{\prime}[T]\supseteq H^{\prime\prime}[T]\supseteq H^{\prime\prime\prime}[T]$ associated to each tree $T\in\mathcal{T}$. In the following let $B$ be an $f$-capped blocking set for $H$ as in Lemma 20.
@@
  • (Construction of $H^{\prime\prime}[T]$) To construct $H^{\prime\prime}[T]$ from $H^{\prime}[T]$, for every pair $(e,e^{\prime})\in B$ with $e,e^{\prime}\in H^{\prime}[T]$, we delete the first edge $e$, and let $H^{\prime\prime}[T]$ be the remaining subgraph.
  • (Construction of $H^{\prime\prime\prime}[T]$) To construct $H^{\prime\prime\prime}[T]$ from $H^{\prime\prime}[T]$, we delete all edges in $T\setminus\texttt{mst}(H^{\prime\prime}[T])$.
@@
Lemma 26.
Every graph $H^{\prime\prime\prime}[T]$ has weighted girth $>k+1$ (deterministically).
@@
Let $C$ be any cycle in $H[T]$ of normalized weight $\leq k+1$. We will argue that $C$ does not survive in $H^{\prime\prime\prime}[T]$. There are two cases, depending on whether $C$ has a heaviest edge $(u,v)\notin T$, or whether all heaviest edges in $C$ are in $T$.
In the first case, suppose that there is a heaviest edge $(u,v)\in C,(u,v)\notin Q$. By the properties of the blocking set $B$ from Lemma 20, there is $(e,e^{\prime})\in B$ with $e,e^{\prime}\in C$ and $e\notin T$. It is not possible for both edges $e,e^{\prime}$ to survive in $H^{\prime\prime}[T]$, since if they are both selected to remain in $H^{\prime}[T]$ then we will choose to delete $e$ when we construct $H^{\prime\prime}[T]$. Thus $C$ does not survive in $H^{\prime\prime}[T]$.
In the second case, suppose that all heaviest edges in $C$ are also in $Q$. Suppose that all edges in $C$ survive in $H^{\prime\prime}[T]$. Then at least one of the heaviest edges $(u,v)\in C$ will not be in $\texttt{mst}(H^{\prime\prime}[T])$ (since otherwise we could exchange it for a lighter edge on $C$). So we will remove $(u,v)$ when we move from $H^{\prime\prime}[T]$ to $H^{\prime\prime\prime}[T]$. ∎
Lemma 27.
Every graph $H^{\prime\prime\prime}[T]$ has expected weight
$\mathbb{E}\left[w(H^{\prime\prime\prime}[T])\right]\geq\Omega\left(\frac{w(H[T])}{f}\right)-w(T).$
Proof.
Every edge in $T$ is included in $H^{\prime\prime}[T]$ deterministically. Every edge $e\in E(H[T])\setminus T$ survives in $H^{\prime\prime}[T]$ iff the following two independent events both occur:
@@
Thus each edge $e\in E(H[T])$ survives in $H^{\prime\prime}[T]$ with probability $\Omega(1/f)$, implying that
Finally, when we move from $H^{\prime\prime}[T]$ to $H^{\prime\prime\prime}[T]$, we can only possibly delete edges in $T$ and so we delete at most $w(T)$ weight, implying the lemma. ∎
@@
As in the probabilistic method, there exists a realization of the subgraphs $H^{\prime\prime\prime}[T]$ that satisfies the expected weight inequality in the previous lemma (for all $T$). Rearranging the bound from Lemma 27, we have:
$w(H^{\prime\prime\prime}[T])$ $\geq\Omega\left(\frac{w(H[T])}{f}\right)-w(T)$
$\implies O(f)\cdot\left(w(H^{\prime\prime\prime}[T])+w(T)\right)$ $\geq w(H[T]).$
Using this, we observe that for all trees $T\in\mathcal{T}$, we have
$\frac{w(H[T])}{w(T)}$ $\leq\frac{O\left(f\right)\cdot\left(w(H^{\prime\prime\prime}[T])+w(T)\right)}{w(T)}$
$\leq O(f)\cdot\frac{w(H^{\prime\prime\prime}[T])}{w(\texttt{mst}(H^{\prime\prime\prime}[T]))}+O(f)$
$=O(f)\cdot\ell(H^{\prime\prime\prime}[T])+O(f)$
$\leq O(f)\cdot\lambda(n,k+1)$ (2)
where the last steps follow from the definition of $\lambda$, the fact that $\lambda(n,k+1)\geq 1$, and the previous lemma establishing that $H^{\prime\prime\prime}[T]$ has weighted girth $>k+1$. Thus, to wrap up the proof, we bound:
@@ 3.4 Improved Lightness Bounds via Multiple Host Trees
A way in which our previous analyses are suboptimal is that they do not acknowledge the possibility that an edge could be hosted by many trees, not just one. As a simple example to introduce the technique, let us observe that the lightness bound from Algorithm 1 improves by a factor of $f$ if we are willing to take a higher lightness competition parameter:
@@
Follow the previous analysis, but note that our guarantee is now that we have a set of trees $\mathcal{T}$ such that (1) every edge is in at most two trees, and (2) for all $(u,v)\in E(G)\setminus Q$, there are $(2+\eta)f+1$ trees in $\mathcal{T}$ that span the $(2+\eta)f+1$-connected component that contains $u,v$. Thus, letting $F$ be a fault set that caused us to add $(u,v)$ to the spanner, there are $\Theta(\eta f)$ of these trees that are disjoint from $F$. We let each of these trees host the edge $(u,v)$. In particular, this means that $(u,v)$ will appear in $\Theta(\eta f)$ many $H[T]$ subgraphs. From there the analysis continues as before, but our final calculation is:
@@ 3.5 Competitive Lightness Upper Bounds in Polynomial Time
The Chernoff bound states that:
@@
The following lemmas will assume that the event from Lemma 30 holds, and hence these lemmas only hold with high probability, rather than deterministically. We next establish correctness:
@@
There are $|F|\leq f$ edges, which are each included in $H^{\prime}$ with probability $1/f$, and so none of the edges in $F$ survive in $H^{\prime}$ with probability at least $1/4$. So we have $P_{e}^{T}\geq 1/4$. Assuming that the event from Lemma 30 holds (relating $P^{T}_{e}$ to $\widehat{P}_{e}^{T}$), we thus have $\widehat{P}_{e}^{T}\geq 1/8$. So $(u,v)$ will be added to the spanner $H$ when the tree $T$ is considered, completing the proof. ∎
@@
Analogous to the previous warmup, it will be helpful to define subgraphs
$H^{\prime\prime\prime}[T]\subseteq H^{\prime\prime}[T]\subseteq H^{\prime}[T]\subseteq H[T]$
as follows. First, $H^{\prime}[T]$ is a random subgraph of $H[T]$, obtained by including $T$ deterministically and then including each edge hosted by $T$ with probability $1/f$. For an edge $(u,v)\in H^{\prime}[T]$, let $H^{\prime}_{(u,v)}[T]$ be the subgraph that contains only $T$ and the edges from $H^{\prime}[T]$ considered strictly before $(u,v)$ in the algorithm. Then we define $H^{\prime\prime}[T]$ as:
@@
Finally, we define
$H^{\prime\prime\prime}[T]:=H^{\prime\prime}[T]\setminus\left(T\setminus\texttt{mst}(H^{\prime\prime}[T])\right).$
We can now bound the expected weight of $H^{\prime\prime\prime}[T]$ in two different ways:
  • On one hand, the definition of $H^{\prime\prime\prime}[T]$ implies that it has weighted girth $>k+1$. The argument is analogous to Lemma 26. Consider a cycle $C$ in $H^{\prime}[T]$ of normalized weight $\leq k+1$. We consider two cases, by whether or not there is a heaviest edge $(u,v)\in C$ that is not also in $T$.
    First suppose that there is a heaviest edge $(u,v)\in C,(u,v)\notin T$, and let $(u,v)$ specifically be the last such edge considered by the algorithm. If all other edges in $C\setminus\{(u,v)\}$ are added to $H^{\prime\prime}[T]$ then then they form a $u\leadsto v$ path, implying that
@@
    which will cause us to not include $(u,v)$ in $H^{\prime\prime}[T]$.
    On the other hand, suppose that all heaviest edges $(u,v)\in C$ are also in $T$. Then $\texttt{mst}(H^{\prime\prime}[T])$ will omit at least one such edge (since otherwise we could exchange it for a lighter edge on $C$), so $C$ will not survive in $H^{\prime\prime\prime}[T]$.
    Putting these together: since $H^{\prime\prime\prime}[T]$ has weighted girth $>k+1$, by definition of $\lambda$, we have
    $\lambda(n,k+1)\geq\ell(H^{\prime\prime\prime}[T])$ $=\frac{w(H^{\prime\prime\prime}[T])}{w(\texttt{mst}(H^{\prime\prime\prime}[T]))}$
    $\geq\frac{w(H^{\prime\prime\prime}[T])}{w(T)}.$
@@
    We then delete at most $w(T)$ additional edge weight when we move from $H^{\prime\prime}[T]$ to $H^{\prime\prime}[T]$, so we have
    $\mathbb{E}\left[w(H^{\prime\prime\prime}[T])\right]\geq\Omega\left(\frac{w(H[T])}{f}-w(T)\right).$
@@
$w(H[T])$ $\leq O\left(f\cdot\mathbb{E}\left[w(H^{\prime\prime\prime}[T])\right]\right)$
@@
The first part of Theorem 13 is now directly implied by Lemmas 31 and 33.
@@ 5 Future Directions: Light Vertex Fault Tolerant Spanners
In addition to the obvious open questions of a) determining the precise bound on the achievable $2f$-competitive lightness, and b) providing simultaneous bounds on both the lightness and the sparsity, an interesting open problem left by this paper is to determine the appropriate competition parameter for light vertex fault tolerant (VFT) spanners. We can define vertex competitive lightness analogously to Definition 8, but with respect to the set of $f$-VFT connectivity preservers rather than $f$-EFT connectivity preservers. We then ask:
Question 35.
What is the smallest function $\mathcal{C}(f)$ such that every $n$-node weighted graph $G$ has an $f$-VFT $k$-spanner with vertex-competitive lightness [^9] It would also be interesting if, instead of a $\lambda$ reduction, one could match the state-of-the art non-faulty size/stretch tradeoff up to a $\texttt{poly}(f)$ factor. That is, the goal would be to obtain stretch $(1+\varepsilon)(2k-1)$ and lightness $O\left(\texttt{poly}(f,\varepsilon^{-1})\cdot n^{1+1/k}\right).$